OpenStudy (anonymous):

125^125+124^124...........1^1 in scientific notation

OpenStudy (anonymous):

\[1.3032773*10^{262} \]\[\sum _{n=1}^{125} n^n= \]13032773189153331238442856175527161721020329077218828118720708 291408647774446025288621114128988466208564442404561288587800234330632046984090740932945287127466402649416618750210255006101848420661996674141915354351661559813795628518919705775046618225800676444866117

OpenStudy (anonymous):

scientific notation, but u did well already :D

OpenStudy (anonymous):

Check out the first line above the Sigma symbol.

OpenStudy (anonymous):

ohhhhh